Title

Article text.

04:40 Collimate MODS1 and MODS2

04:43 Binocular preset to BS9122
  • was not in binocular mode (I decided to send the script to both sides, but didn't tell David), so both scripts failed, but "r" to retry worked fine.
  • acqBinoMODS BS9122.acq
    • 30-sec slit image: mods2r 0001 and mods1r 0003
    • 15-sec field image: saturated on MODS2, but not on MODS1 (only 30k)... ??
    • 10-sec field image also saturated on MODS2 (0003)
  • acqBinoMODS BS9122.acq where the star is a nearby R=14.82 star
    • 30-sec slit image: mods2r 0004 and mods1r 0005 and
    • 15-sec field image mods2r 0005 and mods1r 0006
  • execBinoMODS BS9122slitscan.img
    • track preset which uses detxy dithers to go up/down along the slit. I had intended to keep these within the copointing limit, but I ran the wrong script which did not send any preset and would have sent offsets greater than the copointing limit.
  • execBinoMODS BS9122track.img (mods2r: 7-13) and (mods1r: 8-14)
  updatepointing
  GO
  offsetxy 0 30 abs
  GO
  offsetxy 0 -30 abs
  GO
  offsetxy 0 15 abs
  GO
  offsetxy 0 -15 abs
  GO
  offsetxy 0 40 abs
  GO
  offsetxy 0 -40 abs
  GO
  • Had an offset 0 -30 abs timeout. I hit "r" to retry, probably while the exposure was already in progress so we see a trailed image (mods2 0009).
  • Then, ran into a 'race' condition where it said it was already exposing. Waited a while before "r" but still out of sync...
>> offsetxy 0 30 abs
DONE: offsetxy ABS Slit Offset Completed lbtSide=right
>> GO
STATUS: RED GO Finished 1 of 1 exposures. ACQTAG=MODS2R.20160214051445 EXPSTATUSDONE: RED GO Exposures finished.
>> offsetxy 0 -30 abs
** ERROR: Command 'offsetxy 0 -30 abs' timed out after 30 seconds
** Abort, Retry, or Ignore? > r
** Retrying command 'offsetxy 0 -30 abs'
>> offsetxy 0 -30 abs
DONE: offsetxy ABS Slit Offset Completed lbtSide=right
>> GO
DONE: offsetxy ABS Slit Offset Completed lbtSide=right
>> offsetxy 0 15 abs
DONE: offsetxy ABS Slit Offset Completed lbtSide=right
>> GO
** ERROR: GO cannot start a red exposure - red channel already exposing
** Abort, Retry, or Ignore? > r
** Retrying command 'GO'
>> GO
** ERROR: GO cannot start a red exposure - red channel already exposing
** Abort, Retry, or Ignore? > r (after waiting for a while)
** Retrying command 'GO'
>> GO
STATUS: RED GO Finished 1 of 1 exposures. ACQTAG=MODS2R.20160214051913 EXPSTATUSDONE: RED GO Exposures finished.
>> offsetxy 0 -15 abs
STATUS: RED GO EXPSTATUS=ERASE

05:32 Will try again... we had this problem of an offset timing out when running slitscan script in May 2014, but there the guide star was (if I remember well) close enough to the bottom of the patrol field that this offset would have pushed the probe near the bottom of the field. We thought that might have been a problem, but no... this was a track preset.
  • acqBinoMODS BS9122.acq (seeing is 1.50 and 1.6"...)
  • execBinoMODS BS9122track.img
    • timed out in the same place "0 -30 abs". I aborted the script.
  • acqMODS BS9122.acq --- wrong MODS1
  • acqMODS2 BS9122.acq --- 16 is the slitimage2.fits
  • execMODS2 BS9122track.img --- 0 -30 failed again
>> offsetxy 0 -30 abs
** ERROR: Command 'offsetxy 0 -30 abs' timed out after 30 seconds
** Abort, Retry, or Ignore? > i
** Ignoring error and continuing...
  • do manual offsets: offsetxy 0 15 abs, go (20), offset 0 -30 abs (completed, Rick may need to increase his timeout threshold in execMODS2), go (21), offset 0 -15, go (22), offset 0 35, go (23),
06:28 - determine that 0.5 deg must be added to the current rotator zeropoint.

06:30 - Slew to StoneE1f. Failed, copointing limit. David cancelled left preset. acqMODS2 mods.44.StoneE1f.acq
lbto@obs4:8 % modsAlign_test mods.44.StoneE1f.mms mods2r.20160214.0025.fits mods2r.20160214.0026.fits

MODS2 Offset Command:
   offsetpointing  0.7201 -1.928 -3.566 detxy rel

field image taken after offset applied was 000

06:52 - David restarted PCS after we changed right rotator zeropoint from 27.55 -> 26.95.

07:00 - (will retake slitscan). acqMODS2 mods.44.StoneE1f.acq
modsAlign_test mods.44.StoneE1f.mms mods2r.20160214.0025.fits mods2r.20160214.0026.fits
MODS2 Offset Command:
   offsetpointing  0.7134 -1.908 -3.577 detxy rel

07:08 Send preset again so as to take new set of acq images without the double-offset.
lbto@obs4:13 % modsAlign_test -q mods.44.StoneE1f.mms mods2r.20160214.0030.fits mods2r.20160214.0031.fits
MODS2 Offset Command:
   offsetpointing  0.1088 -4.621 -2.634 detxy rel

07:22 right rotator zeropoint from 26.95 ---> 26.85. David restarted PCS.

07:25 acqMODS2 mods.44.StoneE1f.acq
lbto@obs4:15 % modsAlign_test -q mods.44.StoneE1f.mms mods2r.20160214.0033.fits mods2r.20160214.0034.fits
MODS2 Offset Command:
   offsetpointing  0.0439 -4.937 -2.391 detxy rel

Confirmatory image 35 looks good.

07:33: David switches to track mode red go (36) updatepointing offset 0 15 red go (37) offset 0 -15 red go (38) offset 0 40 red go (39) offset 0 -40 red go (40)
Line of star centroids is 0.181+/-0.001 degrees CCW from the Y axis
Slit is                   0.167+/-0.000 degrees CCW from the Y axis 
0.014+/-0.001 degrees should be added to the current rotator zeropoint 

07:55 Confirmed: the right rotator zeropoint for MODS2 is 26.85 deg

07:56: Take transform data
  • iif_register, /tcs
  • transform_collect
    • /home/lbto/idl/sext/transform/lists/Stone_E_mods_1.list
    • starts with mods2r.20160214.0042.
    • seeing ~ 1.2-1.4" on the guider, but WFS spots ~1. At this low airmass (1.8), DAR is elongating the guide star image.
 Image mods2r.20160214.0042.fits OBJECT - Stone_E_mods_1 On Axis - r_sdss - Imaging IN
 Image mods2r.20160214.0043.fits OBJECT - Stone_E_mods_1 BG - r_sdss - Imaging IN

 Image mods2r.20160214.0044.fits OBJECT - Stone_E_mods_1 GS=81 - r_sdss - Imaging IN (ok)
 Image mods2r.20160214.0045.fits OBJECT - Stone_E_mods_1 GS=87 - r_sdss - Imaging IN (guide probe vignettes on-axis star, but others indicate no motion)
 Image mods2r.20160214.0046.fits OBJECT - Stone_E_mods_1 GS=117 - r_sdss - Imaging IN (on axis star moved right and up)
 Image mods2r.20160214.0047.fits OBJECT - Stone_E_mods_1 GS=72 - r_sdss - Imaging IN (on axis star moved left and up)
 Image mods2r.20160214.0048.fits OBJECT - Stone_E_mods_1 GS=83 - r_sdss - Imaging IN (target under probe shadow)
 Image mods2r.20160214.0049.fits OBJECT - Stone_E_mods_1 GS=104 - r_sdss - Imaging IN (on axis star pretty close to on-axis but a bit low)
 Image mods2r.20160214.0050.fits OBJECT - Stone_E_mods_1 GS=63 - r_sdss - Imaging IN (on axis star pretty close)
 Image mods2r.20160214.0051.fits OBJECT - Stone_E_mods_1 GS=60 - r_sdss - Imaging IN (on axis star way up)
 Image mods2r.20160214.0052.fits OBJECT - Stone_E_mods_1 GS=47 - r_sdss - Imaging IN (on axis star close to on axis)
 Image mods2r.20160214.0053.fits OBJECT - Stone_E_mods_1 GS=39 - r_sdss - Imaging IN (ditto)
 Image mods2r.20160214.0054.fits OBJECT - Stone_E_mods_1 GS=99 - r_sdss - Imaging IN (ditto)
 Image mods2r.20160214.0055.fits OBJECT - Stone_E_mods_1 GS=90 - r_sdss - Imaging IN (on axis star far away)
(on guide star 91, preset failed - star too far off axis ... couldn't move probe to position (out of patrol field) (SFP) (x,y): 74.2222, -212.862 mm).
 Image mods2r.20160214.0056.fits OBJECT - Stone_E_mods_1 GS=102 - r_sdss - Imaging IN (on axis star close)
 Image mods2r.20160214.0057.fits OBJECT - Stone_E_mods_1 GS=77 - r_sdss - Imaging IN (on axis star close)
 Image mods2r.20160214.0058.fits OBJECT - Stone_E_mods_1 GS=42 - r_sdss - Imaging IN (on axis star close)

Summary: Many guide stars were incorrectly selected and the science field shifted as a result. These are noted above, and, in all, we probably only have about 9 useful images. Is Stone E just too dense for transform data collection? But, it is used for the LUCIs, so GCS must not be choosing the incorrect guide star for the LUCI AGw/W's.

08:46 Reconfiguring for LUCI2

=====================================

UT 09:07 - moving to WAS 49 #087 first science exposure - guider 0.87 WFS

UT 09:20 - starting science integration

UT 09:22 - error committing instrument setup - cannot find grating name -

UT 09:23 - Putting the G200 in to test - OK

UT 09:24 - Putting G210 in - could not find grating name

UT 09:26 - Backup - switch to G200 HK CWL 1.95 - switch is OK - script vanished from scheduler

UT 09:35 - Re-running this at G200HK CWL 1.95 - made new script - all ok so far

UT 09:39 - Taking data - first exposure #95, last 104

UT 10:06 - moving to Telluric Star HIP 62194

UT 10:11 - tilde issue with RTD2 - filename comes up as luci2.2016214~1

UT 10:13 - RTD2 froze after clicking RESET ALL to load images properly

UT 10:13 - Observer GUI frozen

UT 10:18 - taking spectra of std

UT 10:19 - Error commit - went to 1micron for CWL - not sure why - script looks correct - checked in scriptor

UT 10:21 - half the script vanished

UT 10:23 - reset script - skipped to start of spectroscopic integration

UT 10:24 - same error

UT 10:26 - re-acquiring, failed to find guidestar

UT 10:26 - checking pointing

UT 10:29 - sending Telluric acquisition again

UT 10:38 - defaulting to 1.0micron

UT 10:38 - going to 1.93 CWL to see if happy

UT 10:40 - trying 1.93 again - failed again

UT 10:41 - running initialize all except MOS

UT 10:42 - doing manually, except, its defaulted to NORMAL not integrated - but didn't save 20 images 110-129 are missing

UT 10:44 - taking an integrated science exposure

UT 10:45 - manually offsetting 0,+20 in detxy with telsvc

UT 10:50 - doing manually 0,-20.0, taking image manually

UT 10:52 - USE 132,133 for Telluric

UT 10:52 - Moving back to WAS49 field to do Ks imaging - seeing really good

UT 10:57 - checking pointing

UT 11:02 - seeing bloomed? was 0.6" with telluric, now its 1.1-1.3" for imaging

UT 11:15 - back to sub-arcsecond seeing

UT 11:16 - Imaging script part 2 (same field - different offsets)

UT 11:23 - FM4 slipped again

UT 11:32 - moving to FS33 std

UT 11:35 - Airmass 1.041

UT 11:40 - Completed WAS49 Program - last science exposure #157

UT 11:41 - putting blinds in filter wheel

UT 11:41 - Reconfiguring to LBCs for WRS testing
  • John is installing new WRS code from Marco
  • We intend to measure focus offsets and to try WRS in fast mode

UT 12:04 - David slewed to Persson star, BS9162 (15:59:13, 47:36:41)
  • dowrs, /blueonly --- permissions problem with a new procedure, cartop.pro, caused dowrs to crash. John changed permissions.
  • dowrs,/blueonly
    • converged, however there is a lot of spherical and the pupil images (065536) look like rings.

UT 12:16 - David clearing active optics on SX. And he is giving a -0.3 mm offset to the Z global offset of -0.3 mm.
  • dowrs, /blueonly ---
    • 121835 has too much spherical, yet dowrs sent +105nm z11.
    • 122002 looks better, though.
    • 122125. After this, it declared convergence, but there is some coma. Very little coma was sent.
  • dowrs, /blueonly again ... we're thinking we may need to turn back on the interactive coma. "S" to stop this. Edited WRS_BLUE.par and WRS_RED.par to turn INTERACTIVE back to 1.
  • dowrs, /blueonly "S" because I realized I had not connected to TCS. Clicked "Connect to LBT" on LBC UI. dowrs was still sending Zernikes, however. TCSSendWavefront does not check whether the LBC is connected. The OB does, however - not having TCS connected was why we thought we needed the -0.3 mm offset to the global offset.
  • dowrs, /blueonly -- dowrs was sending very small coma corrections. When prompted, the final time, I exaggerated my choices of the centers of the pupil circumference and inner obscuration, dowrs gave z7 = -193 nm and then declared it had converged. I'm not sure it needed this 'kick', as I thought that the choice of center didn't have to be so exact - this may have been a coincidence.

UT 12:40 - Superfoc sequence started.

UT <~12:50 - David entered Zglobal offset on DX of -1.3mm (-1.00 mm from collim model), and also restored Zglobal on SX to -0.3mm.
  • dowrs, /redonly --- converged but did not seem to remove coma (very small z7/z8 corrections sent)
  • repeating dowrs, /redonly...

UT 13:05 - Superfoc sequence started.

UT 13:08 - John noted that the log file, WRS_report.dat that is written in /home/lbceng/WRS/src/ is not being updated today. I am running dowrs from my lbto account. The log file is being saved to /home/lbto tonight (it is attached here, also).

UT 13:20 - trying the fast mode: I edited WRS_BLUE.par to set the FAST flag to 1, then ran dowrs, /blueonly. Looks like it may be too bright (we're into twilight):


RUNNING WRS IN FAST MODE
WARNING NO PUPIL DETECTED TRYING AUTOMATIC NOISE THRESHOLD ADJUSTMENT ...
Threshold adjustment:       2.70000 sigma
Threshold adjustment:       2.43000 sigma
Threshold adjustment:       2.18700 sigma
Threshold adjustment:       1.96830 sigma
Threshold adjustment:       1.77147 sigma
Threshold adjustment:       1.59432 sigma
Detection threshold is  too low. WRS cannot proceed. Check configuration fiels or the image.

UT 13:27 closing the enclosure and setting up for LUCI2 calibrations.

UT 13:34 - starting G200 HKspec cals #158-183.

UT 13:42 - DONE

UT 13:44 - LUCI-2 blind,blind, 0.5" in FPU, flex off, cal out, N3.75

==============================

-- %USERSIG{OlgaKuhn - 2016-02-14}%

Comments

 
Topic revision: r7 - 14 Feb 2016, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback